OpenStudy (anonymous):

what is the diagonal of a square when the sides are 16 inches?

OpenStudy (goldphenoix):

 Do you have any idea on how to solve it?

OpenStudy (anonymous):

No. This is the question i'm trying to solve and I just can't get it.

OpenStudy (anonymous):

I'm doing: SIMILAR POLYGONS THEOREM ABOUT 45-45-90 RIGHT TRIANGLES

OpenStudy (goldphenoix):

^I forgot how to do that. But do you know the pythagoream theorm? That can help you answer the problem.

OpenStudy (goldphenoix):

The pythagoream theorm looks like: \[\large \large a^2 + b^2 = c^2\] A and B represent the legs, while C represent the hypotenuse. So what do you think you have to do?

OpenStudy (anonymous):

It doesn't give me the leg only the hypotenuse.

OpenStudy (goldphenoix):

Oh really?  Do you see it now?

OpenStudy (goldphenoix):

I have to go soon, but you see the triangle, right? The 2 leg (a,b) is 16. So the equation would look like: \[\large \large 16^2 + 16^2 = c^2\] Solve your equation, and c would be the diagonial of the square. :)

OpenStudy (anonymous):

Yes, I see it. Thanks.